# A Comprehensive Survey of Synthetic Tabular Data Generation

### v2

@@ Abstract
Tabular data remains one of the most prevalent and critical data formats across a wide range of real-world applications. However, its effective utilization in machine learning (ML) is frequently constrained by challenges such as data scarcity, privacy concerns, and class imbalance. Synthetic data generation has emerged as a promising solution, leveraging generative models to learn the underlying distribution of real datasets and produce high-fidelity, privacy-preserving synthetic samples. A variety of generative paradigms have been explored for this task, including energy-based models (EBMs), variational autoencoders (VAEs), generative adversarial networks (GANs), large language models (LLMs), and diffusion models. While several existing surveys have investigated synthetic tabular data generation, the majority tend to concentrate on narrow subdomains or specific generative approaches, such as focusing exclusively on GANs, diffusion models, or privacy-preserving techniques. This narrow scope often results in fragmented perspectives, lacking a comprehensive, comparative synthesis that bridges diverse methodological paradigms. In particular, recent progress driven by emerging generative models, such as LLMs and diffusion-based techniques, has not been adequately covered. This gap in the literature impedes a holistic understanding of the field’s current landscape, the interplay between different approaches, and key open challenges. To bridge this gap, this survey offers a unified and systematic review of synthetic tabular data generation. Our contributions are threefold: (1)we present a comprehensive taxonomy that organizes existing methods into three distinct categories-traditional generation methods, diffusion model methods, and LLM-based methods-and provide an in-depth comparative analysis for each approach; (2)we provide a detailed overview of the entire pipeline for synthetic tabular data generation, covering synthetic tabular data generation, post processing, and evaluation; (3)we identify key challenges in the field, explore real-world application domains, and outline open research questions and future directions to guide further progress in this rapidly evolving area; (4)We have consolidated the relevant paper, source code, and datasets into a publicly accessible code repository [^1] GitHub repository.
@@ I Introduction
Tabular tabular data is a fundamental data format in real-world applications, playing a crucial role in domains such as healthcare [1], finance [2], education [3], transportation [4], and psychology [5]. In the face of strict data privacy regulations such as the General Data Protection Regulation (GDPR) [6] and California Consumer Privacy Act (CCPA) [7], access to and sharing of data are severely restricted. Furthermore, the inherent challenges of missing values and imbalanced class distributions in tabular data further limit its availability in practical applications. Fortunately, generative models offer a promising solution to these challenges by learning the distribution of real data and generating synthetic samples that resemble the original data while mitigating privacy concerns and data limitations.
@@
Generative models such as Energy-based Models (EBMs) [8], Variational Autoencoders (VAEs) [9], Generative Adversarial Networks (GANs) [10], Normalizing Flows [11], Large Language Models (LLMs) [12], and Diffusion Models [13] have been employed to synthesize tabular data with a distribution similar to the original data. Although existing generative methods have achieved remarkable performance on homogeneous data such as text and images, they often struggle to generate high-quality tabular data due to the inherent heterogeneity and complex distributional characteristics of tabular datasets. This challenge has attracted widespread interest and led to extensive research in the field. Such as TabDDPM [14], TABDIFF [15] solve heterogeneity problems through multimodal diffusion. CLLM [16] leverages the internal knowledge of LLMs to improve generation quality in low-data scenarios. DP-LLMTGen [17], DPGAN [18], PATE-GAN [19] enhance the privacy protection of both generative models and their outputs by incorporating differentially private strategies.
@@
Numerous survey papers have been published to summarize related work from various perspectives. However, existing surveys have several limitations: (1) Some existing works are limited to specific domains or methods and lack a comprehensive overview of synthetic tabular data generation. Such as some surveys [20, 21, 22] focus on the use of GANs for generating medical tabular data, some surveys [23, 24] focus on class imbalance, privacy, and fairness in synthetic tabular data generation. (2) Existing work primarily focuses on the generation and evaluation of synthetic tabular data [25, 26], while overlooking the importance of post-processing after data generation. Synthetic data may match the original distribution but still violate commonsense or logical rules (e.g. negative age values or children marked as married). Post-processing is therefore necessary to ensure semantic correctness and data validity. (3) The latest advancements in LLMs and diffusion models for synthetic tabular data generation have not yet been thoroughly explored [27]. LLMs and diffusion models offer advantages over traditional generative methods—LLMs for their ability to model semantic structure, and diffusion models for their stability and representation capabilities—but current surveys lack a thorough investigation of these methods.
@@
In this survey, we present a comprehensive review of synthetic tabular data generation. Specifically, we first introduce the definition of synthetic tabular data generation and the key challenges it faces. We then categorize existing methods into three main classes—traditional generation methods, diffusion model methods, and LLM-based methods—and provide a detailed analysis of the strengths and limitations of each category. Furthermore, we summarize and classify the post-processing techniques into two primary categories: sample enhancement and label enhancement, highlighting their importance in ensuring data quality and semantic consistency. We also review evaluation methodologies from the perspectives of data availability and privacy, reflecting the dual objectives of synthetic tabular data. Finally, we discuss real-world applications of synthetic tabular data and outline potential directions for future research.
The main contributions of this work are as follows:
  • We categorize existing methods into three major classes: traditional generation methods, diffusion model methods, and LLMs-based methods. A comprehensive review of each category is conducted to provide future researchers with a structured foundation for further exploration and development in this field.
  • In this work, we present an overview of the entire workflow of synthetic tabular data generation, emphasizing the importance of post-processing for ensuring the quality and validity of the synthetic data. We also provide a comprehensive summary of existing post-processing techniques to support and guide future research in this area.
  • We provide a detailed survey of the latest advancements in synthetic tabular data generation and summarize its key challenges, practical applications, and future opportunities.
  • We summarize widely used datasets for synthetic tabular data generation, detailing their sources, associated tasks, and application domains to support future research.
@@ II-A Problem Definition
In this section, we formally define the problem settings for synthetic tabular data generation: rather than constructing a dataset from scratch, synthetic tabular data generation focuses on learning the statistical distribution of an existing tabular dataset $D_{e}$ and producing a synthetic tabular dataset $D_{s}$ that closely approximates it. Consequently, the overall generation task can be formulated as:
@@
where $p_{\theta}$ represents the generation model and $\mathcal{D}_{e}=\left\{(x_{e}^{i},y_{e}^{i})\right\}_{i=1}^{N_{l}}$ denote a tabular dataset consisting of $N_{l}$ samples, where $x_{e}^{i}$ indicates all feature values of $i$-th row, and $y_{e}^{i}$ represents its corresponding label. $D_{e}$ and $D_{s}$ follow the same structured format defined by a set of feature names $F=\{f_{j}\}_{j=1}^{d}$, where $f_{j}$ represents the name of the $j$-th feature, and $d$ denotes the total number of features.
Since $D_{s}$ often contains mislabeled and violate common human knowledge or domain-specific constraints samples, post-processing techniques are used to further improve the quality of the $D_{s}$ and ensure alignment with human knowledge, which include Sample Enhancement and Label Enhancement, see Section IV for more details. Moreover, a comprehensive evaluation of data availability and privacy is also important to generate high-quality $D_{s}$, in Section V, we summarize the evaluation metrics for data availability, including ML efficiency, fidelity, and alignment, as well as the metrics for privacy, such as Distance to Closest Record (DCR), Attribute Inference Attack, and Membership Inference Attack. These steps help ensure that $D_{s}$ can be effectively used in real-world applications while maintaining its privacy and reliability. The complete pipeline of synthetic tabular data generation is illustrated in Figure 1.
@@ II-B Challenges for Synthetic Tabular Data Generative Models
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=8.0em,font=,, where level=3text width=9.5em,font=,, where level=4text width=12em,font=,, [Synthetic Tabular Data Generation (§III),ver [Traditional Generation
@@
AutoDiff [57], RelDDPM [58], leaf2, text width=32.2em] ] [Medical, text width=8em [e.g.,  TabDDPM-HER [59], DPM-HER [60], FLEXGEN-EHR [61], leaf2, text width=32.2em] ] [Financial, text width=8em [e.g.,  EntTabDiff [62], FinDiff [63], Imb-FinDiff [64], leaf2, text width=32.2em] ] ] [Score-based
@@
Synthetic tabular data is generative unlike synthetic text and image data, which consist of uniform basic units (e.g., pixels or tokens) that allow the generative model $p_{\theta}$ to learn distributions in a homogeneous space, tabular data is inherently heterogeneous and structurally diverse, posing substantial challenges for both effective modeling and the generation of high-quality synthetic data. In the following, we present the key challenges of synthetic tabular data generation from three perspectives: data quality, inherent characteristics of tabular data, and complex and diverse data distributions.
@@ II-B1 Data Quality
  • Limited Data Size. Tabular data, particularly in specialized domains such as healthcare, finance, and agriculture, is often challenging to collect and annotate. As a result, most datasets are limited in size, which significantly restricts their availability. This data scarcity poses challenges for training generative models and ensuring the reliability of $D_{s}$.
  • Imbalanced Class Distribution. Tabular data frequently exhibits severe class imbalance, where the data size of certain categories is significantly lower. For instance, in medical datasets, the number of patients is often much lower than that of healthy individuals. Similarly, in manufacturing, defective products make up only a small fraction of the data. Such imbalance makes it difficult for generative models to learn accurate representations of minority classes, often resulting in biased synthetic outputs.
  • Missing Values. Missing values are common in tabular datasets due to privacy constraints, human error, or the optional nature of certain attributes. For example, in electronic health records, not every diagnostic test is performed for every patient, leading to incomplete entries. These missing values increase modeling complexity, as generative models must learn underlying distributions while handling incomplete observations.
@@ II-B2 Inherent Characteristics
  • Heterogeneity. Unlike homogeneous data types (e.g., pixels or tokens), tabular data contains a mix of categorical and numerical features, each with distinct statistical properties. This structural heterogeneity makes it difficult for conventional generative models to effectively model the joint distribution of such mixed data.
  • Complex and Sparse Feature Dependencies. Tabular data often contains complex inter-feature relationships, including both homogeneous (e.g., numerical-numerical) and heterogeneous (e.g., categorical-numerical) dependencies. However, these relationships are typically sparse—only certain feature pairs may be correlated. For instance, age and education level might be interrelated, whereas others remain independent. Accurately capturing these sparse and complex dependencies is crucial to preserving realism in synthetic data $D_{s}$.
  • Mixed-Type Single Features. In some cases, individual features within a tabular dataset may contain mixed data types, such as numerical values and categorical labels. This further complicates modeling, as it challenges traditional encoding strategies and increases the difficulty of learning coherent representations within a single column.
II-B3 Complex and Diverse data Distributions
The complex and diverse distributional patterns in tabular data bring additional challenges to the generation of high-quality synthetic data $D_{s}$.
  • Non-Gaussian Distribution of Numerical Features. Numerical attributes in tabular data are often different from Gaussian distributions, exhibiting skewness, heavy tails, or multimodal characteristics. This poses a challenge for existing generative models to learn such complex distributions.
  • Imbalanced Categories in Categorical Features. Categorical columns often exhibit Imbalanced distributions, where certain categories dominate the feature. Generative models may overfit these dominant classes, leading to poor representation of minority classes and reducing the diversity of $D_{s}$.
  • Marginal Effects of Columns. The same values may have different meanings depending on the feature context. For example, the value “10” could represent an age, a rating, or a quantity, depending on the feature. Maintaining these semantic distinctions is essential to preserve the interpretability and correctness of $D_{s}$.
@@ III Synthetic Tabular Data Generation Methods
In this section, we based on the employed techniques systematically categorize synthetic tabular data generation methods into traditional generation methods, diffusion model methods, and LLM-based methods, as shown in Figure LABEL:figure:method. Additionally, we analyze the advantages and limitations of each approach to provide a comprehensive and reliable reference for future research in synthetic tabular data generation.
@@ III-A Traditional Generation Methods
GAN-based Method medGAN [40] 2017 MLHC Traditional
TGAN [39] 2018 arXiv Traditional
CrGAN-Cnet [42] 2018 arXiv Traditional
TableGAN [43] 2018 PVLDB Traditional
IT-GAN [44] 2021 NeurIPS Traditional
FCT-GAN [46] 2022 arXiv Traditional
TabFairGAN [47] 2022 MAKE Traditional
VT-GAN [45] 2023 arXiv Traditional
CasTGAN [41] 2024 IEEE Access Traditional
cGANS [48] 2018 EXPERT SYST APPL Conditional
CTGAN [38] 2019 NeurIPS Conditional
cWGAN [51] 2020 EXPERT SYST APPL Conditional
CTAB-GAN [49] 2021 ACML Conditional
OCT-GAN [50] 2021 WWW Conditional
RC-TGAN [52] 2023 ICASSP Conditional
DPGAN [18] 2018 arXiv Differential Privacy
PATE-GAN [19] 2019 ICLR Differential Privacy
RDP-CGAN [53] 2022 Information Sciences Differential Privacy
CTAB-GAN+ [54] 2024 Frontiers Big Data Differential Privacy
@@
In the early stages of synthetic tabular data generation, researchers primarily relied on traditional methods such as Machine Learning Methods, VAE-based methods, and GAN-based methods as shown in Figure 3. Among them, GAN-based methods are further categorized into three types based on their application scenarios: Traditional GANs, Conditional GANs, and Differentially Private GANs, as summarized in Table I.

### v3

@@ Abstract
Tabular data is one of the most prevalent and important data formats in real-world applications such as healthcare, finance, and education. However, its effective use in machine learning is often constrained by data scarcity, privacy concerns, and class imbalance. Synthetic tabular data generation has emerged as a powerful solution, leveraging generative models to learn underlying data distributions and produce realistic, privacy-preserving samples. Although this area has seen growing attention, most existing surveys focus narrowly on specific methods (e.g., GANs or privacy-enhancing techniques), lacking a unified and comprehensive view that integrates recent advances such as diffusion models and large language models (LLMs).
In this survey, we present a structured and in-depth review of synthetic tabular data generation methods. Specifically, the survey is organized into three core components: (1) Background, which covers the overall generation pipeline, including problem definitions, synthetic tabular data generation methods, post processing, and evaluation; (2) Generation Methods, where we categorize existing approaches into traditional generation methods, diffusion model methods, and LLM-based methods, and compare them in terms of architecture, generation quality, and applicability; and (3) Applications and Challenges, which summarizes practical use cases, highlights common datasets, and discusses open challenges such as heterogeneity, data fidelity, and privacy protection.
This survey aims to provide researchers and practitioners with a holistic understanding of the field and to highlight key directions for future work in synthetic tabular data generation.
@@ I Introduction
Tabular data is a fundamental data format in real-world applications, playing a crucial role in domains such as healthcare [1], finance [2], education [3], transportation [4], and psychology [5]. Due to data privacy regulations such as the General Data Protection Regulation (GDPR) [6] and California Consumer Privacy Act (CCPA) [7], the data availability are restricted. Furthermore, the inherent challenges of missing values and imbalanced class distributions in tabular data further limit its availability in practical applications. Fortunately, generative models offer a promising solution to these challenges by learning the distribution of tabular data and generating synthetic samples that resemble the tabular data while mitigating privacy concerns and data limitations.
@@
Generative models such as Energy-based Models (EBMs) [8], Variational Autoencoders (VAEs) [9], Generative Adversarial Networks (GANs) [10], Normalizing Flows [11], Diffusion Models [12], and Large Language Models (LLMs) [13] have been employed to synthesize tabular data with a distribution similar to the real data. Although existing generative methods have achieved remarkable performance on homogeneous data such as text and images, they often struggle to generate high-quality tabular data due to the inherent heterogeneity and complex distribution of tabular data. This challenge has attracted widespread interest and led to extensive research in this field. Such as TabDDPM [14], TABDIFF [15] solve heterogeneity problems through multimodal diffusion. CLLM [16] leverages the internal knowledge of LLMs to improve generation quality in limited data scenarios. DP-LLMTGen [17], DPGAN [18], and PATE-GAN [19] enhance the privacy protection of both generative models and their outputs by incorporating differentially private strategies.
@@
Although several existing studies have provided summaries of methods for tabular data generation, existing surveys have several limitations: (1) Some works are limited to specific domains or methods and lack a comprehensive overview of synthetic tabular data generation, such as, some surveys [20, 21, 22] focus on the GANs for generating medical tabular data, while some surveys [23, 24] focus on class imbalance, privacy, and fairness in synthetic tabular data generation. (2) Some works primarily focuses on the generation and evaluation of synthetic tabular data [25, 26], while overlooking the importance of post-processing after data generation. Synthetic data may match the original distribution but still violate commonsense or logical rules (e.g., negative age values or children marked as married). Therefore post-processing is necessary to ensure semantic correctness and data validity. (3) The latest advancements in diffusion models and LLMs for synthetic tabular data generation have not yet been thoroughly explored [27]. Diffusion models and LLMs offer advantages over traditional generative methods (e.g., diffusion models for their stability and representation capabilities, and LLMs for their ability to model semantic structure), but current surveys lack a thorough investigation of these methods.
@@
In this survey, we present a comprehensive review of synthetic tabular data generation, and the pipeline of synthetic tabular data generation is as shown in Figure 1. We begin by formally defining the task of synthetic tabular data generation in order to clarify its objectives and the challenges. Then, existing approaches are categorized into three categories: traditional generation methods, diffusion model methods, and LLM-based methods, and each category was analyzed in terms of its strengths and limitations. In addition, we summarize and classify the post-processing techniques into two primary categories: sample enhancement and label enhancement, both of which play a critical role in improving data quality and semantic validity. Evaluation methodologies are reviewed from the perspectives of data availability and privacy protection, reflecting the dual objectives of synthetic tabular data. Finally, we discuss the real-world applications and outline promising directions for future research.
The main contributions of this survey are summarized as follows:
  • A deep comparative analysis across paradigms: Building on an established three-fold taxonomy, we conduct a detailed, side-by-side comparison of traditional methods, diffusion models, and LLM-based approaches across multiple dimensions, including data type compatibility, sample fidelity, privacy resilience, and downstream performance.
  • The first systematic synthesis of post-processing strategies: We introduce a novel taxonomy of post-processing methods for synthetic tabular data, distinguishing between sample enhancement and label enhancement techniques. This perspective bridges the gap between data generation and real-world deployment quality requirements.
  • An integrated evaluation framework: We unify existing evaluation practices by organizing them into two principal dimensions, data availability (e.g., ML efficiency, fidelity, alignment) and privacy protection (e.g., membership inference attack, DCR), providing standardized guidance for future benchmarking.
  • A curated repository to promote reproducibility: We compile and release a comprehensive collection of synthetic tabular data generation resources, including representative papers, open-source implementations, datasets, and evaluation tools, available at: https://github.com/ruxueshi/Awesome-Comprehensive-Survey-of-Synthetic-Tabular-Data-Generation.git.
@@ II Definition
To clearly describe the task of synthetic tabular data generation and its associated challenges, this section first provides a formal definition of the overall generation process. Then, we introduce the key challenges from three perspectives: data quality requirements, the inherent characteristics of tabular data, and the complexity of data distributions.
@@ II-A Problem Definition
Synthetic tabular data generation focuses on learning the statistical distribution of an tabular dataset $D_{e}$ and producing a synthetic tabular dataset $D_{s}$ that closely approximates it. Consequently, the overall generation task can be formulated as:
@@
where $p_{\theta}$ represents the generation model and $\mathcal{D}_{e}=\left\{(x_{e}^{i},y_{e}^{i})\right\}_{i=1}^{N_{l}}$ denotes a tabular dataset consisting of $N_{l}$ samples, where $x_{e}^{i}$ indicates all feature values of $i$-th row, and $y_{e}^{i}$ represents its corresponding label. $D_{e}$ and $D_{s}$ follow the same structured format defined by a set of feature names $F=\{f_{j}\}_{j=1}^{d}$, where $f_{j}$ represents the $j$-th feature name, and $d$ denotes the total number of features.
Since $D_{s}$ often contains mislabeled and violate common human knowledge or domain-specific constraints samples, post-processing techniques are used to further improve the quality of the $D_{s}$ and ensure the alignment with human knowledge, which include Sample Enhancement [16] and Label Enhancement [28], see Section IV for more details. Moreover, a comprehensive evaluation of data availability and privacy protection is also important to generate high-quality $D_{s}$, in Section V, we summarize the evaluation metrics for data availability, including ML efficiency [29], fidelity [30, 31], and alignment [32], as well as the metrics for privacy protection [32], such as Distance to Closest Record (DCR), Attribute Inference Attack, and Membership Inference Attack. These steps ensure that $D_{s}$ can be effectively used in real-world applications while maintaining its privacy protection and data availability.
@@ II-B Challenges for Synthetic Tabular Data Generative Models
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=8.0em,font=,, where level=3text width=9.5em,font=,, where level=4text width=12em,font=,, [Synthetic Tabular Data Generation,ver [Traditional Generation
@@
AutoDiff [62], RelDDPM [63], leaf2, text width=32.2em] ] [Domain-specific, text width=8em [e.g.,  TabDDPM-HER [64], FinDiff [65], DPM-HER [66], Imb-FinDiff [67],
FLEXGEN-EHR [68], EntTabDiff [69], leaf2, text width=32.2em] ] ] [Score-based
@@
Synthetic tabular data is more challenges unlike synthetic text and image data, which consist of uniform basic units (e.g., pixels or tokens) that allow the generative model $p_{\theta}$ to learn distributions in a homogeneous space, tabular data is inherently heterogeneous and structurally diverse, posing substantial challenges for both effective modeling and the generation of high-quality synthetic data. In the following, we present the key challenges of synthetic tabular data generation from three perspectives: data quality, inherent characteristics of tabular data, and complex data distributions.
@@ II-B1 Data Quality
  • Limited Data Size. Tabular data, particularly in specialized domains such as healthcare, finance, and agriculture, is often challenging to collect and annotate. As a result, most datasets are limited in size, which significantly restricts their availability. This data scarcity poses challenges for training generative models.
  • Imbalanced Class Distribution. Tabular data frequently exhibits severe class imbalance, where the data size of certain categories is significantly lower. For instance, in medical datasets, the number of patients is often much lower than that of healthy individuals. Similarly, in manufacturing, defective products make up only a small fraction of the data. Such an imbalance makes it difficult for generative models to learn the representations of minority classes, often resulting in biased synthetic outputs.
  • Missing Values. Missing values are common in tabular datasets due to privacy protection constraints, human error, or the optional nature of certain attributes. For example, in electronic health records, not every diagnostic test is performed for every patient, leading to incomplete entries. These missing values increase modeling complexity, as generative models must learn underlying distributions while handling incomplete observations.
@@ II-B2 Inherent Characteristics
  • Heterogeneity. Unlike homogeneous data types (e.g., pixels or tokens), tabular data contains a mix of categorical and numerical columns, each with distinct statistical properties. This heterogeneity makes it difficult for conventional generative models to model the joint distribution of such mixed data effectively.
  • Complex and Sparse Inter-column Dependencies. Tabular data often contains complex inter-column dependencies, including both homogeneous (e.g., numerical-numerical) and heterogeneous (e.g., categorical-numerical) dependencies. However, these relationships are typically sparse—only certain column pairs may be correlated. For instance, age and education level might be interrelated, whereas others remain independent. Accurately capturing these sparse and complex dependencies is crucial to preserving realism in synthetic data $D_{s}$.
  • Mixed-Type Single Column. In some cases, a column of tabular data may contain both numerical and categorical data. This further complicates modeling, as it challenges traditional encoding strategies and increases the difficulty of learning coherent representations within a single column.
II-B3 Complex data Distributions
The complex distribution in tabular data bring additional challenges to the generation of high-quality synthetic data $D_{s}$.
  • Non-Gaussian Distribution of Numerical Columns. Numerical columns in tabular data are often different from Gaussian distributions, exhibiting skewness, heavy tails, or multimodal characteristics. This poses a challenge for existing generative models to learn such complex distributions.
  • Imbalanced Categories in Categorical Columns. Categorical columns often exhibit imbalanced distributions, where certain categories dominate the feature. Generative models may overfit these dominant classes, leading to poor representation of minority classes and reducing the diversity of $D_{s}$.
  • Marginal Effects of Columns. The same values may have different meanings depending on the column name. For example, the value “10” could represent age in age column, or rating in rate column, depending on the which column it belongs to. Maintaining these semantic distinctions is essential to preserving the interpretability and correctness of $D_{s}$.
@@ III Synthetic Tabular Data Generation Methods
In response to the emergence of various generation approaches, we propose a novel taxonomy that categorizes existing methods into three categories: Traditional Generation Methods, Diffusion Model Methods, and LLM-based Methods. For each category, we introduce representative methods, analyze their respective strengths and limitations, and provide a comprehensive and reliable support for future research in this field.
@@ III-A Traditional Generation Methods
GAN-based Method medGAN [45] 2017 MLHC Traditional GAN
TGAN [44] 2018 arXiv Traditional GAN
CrGAN-Cnet [47] 2018 arXiv Traditional GAN
TableGAN [48] 2018 PVLDB Traditional GAN
IT-GAN [49] 2021 NeurIPS Traditional GAN
FCT-GAN [51] 2022 arXiv Traditional GAN
TabFairGAN [52] 2022 MAKE Traditional GAN
VT-GAN [50] 2023 arXiv Traditional GAN
CasTGAN [46] 2024 IEEE Access Traditional GAN
cGANS [53] 2018 EXPERT SYST APPL Conditional GAN
CTGAN [43] 2019 NeurIPS Conditional GAN
cWGAN [56] 2020 EXPERT SYST APPL Conditional GAN
CTAB-GAN [54] 2021 ACML Conditional GAN
OCT-GAN [55] 2021 WWW Conditional GAN
RC-TGAN [57] 2023 ICASSP Conditional GAN
DPGAN [18] 2018 arXiv Differential Privacy GAN
PATE-GAN [19] 2019 ICLR Differential Privacy GAN
RDP-CGAN [58] 2022 Information Sciences Differential Privacy GAN
CTAB-GAN+ [59] 2024 Frontiers Big Data Differential Privacy GAN
@@
In the early stages of synthetic tabular data generation, traditional methods such as machine learning methods, VAE-based methods, and GAN-based methods, as shown in Figure 3 are the main approaches for synthetic tabular data generation. Among them, GAN-based methods are further categorized into three categories based on their application scenarios: Traditional GANs, Conditional GANs, and Differentially Private GANs, as summarized in Table I.